MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2008/04/11
OBSERVER: Fukui Paul
CONDITION: clear -> cloudy -> clear
COMMENT: D3563(300sec), D3565(60sec) are OK. We observed gb2 as the same frequency of gb5 and gb9 for MOA-75.
Compressor gas pressure: 1.9 MPa
Expected Observation Time: START 19:17 END 06:22
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 12441-12476 | 36 | 221 | 13260 |
LMC | 2700-2704 | 5 | 25 | 7500 |
SMC | 634-634 | 1 | 2 | 600 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 42 | 248 | 21360 (05h56m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
8 | 32 | 8 | 17 | 28 | 7 | 8 | 8 | 28 | 7 | 7 | 7 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
7 | 7 | 5 | 6 | 5 | 5 | 5 | 6 | 5 | 5 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 1 | 1 | 1 | 5 | 5 | 4 | 1 | 1 | 1 |
Field13 | Field14 | Field15 | Field16 |
1 | 1 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
18:20 | observing | clear |
19:00 | observing | clear |
20:00 | observing | clear |
21:00 | observing | clear |
22:00 | observing | clear |
22:30 | waiting | cloudy |
23:00 | waiting | partly cloudy |
00:00 | observing | clear |
01:00 | observing | clear |
02:00 | observing | clear |
03:00 | waiting | cloudy |
03:40 | observing | partly cloudy |
05:00 | observing | clear |
06:20 | observing | clear |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day